Please please help! This is a natural logorithm ? I got half, but now I am stumped, and I need it for tomorrow!!! Here it is:
Suppose that the function P=11+44 ln x represents the percentage of inbound e-mail in the U.S. that is considered spam, where x is the number of years after 2002.
Carry all calculations to six decimals on each intermediate step when necessary.
a) Use this model to approximate the percentage of spam in the year 2006 to the nearest tenth of a percent.
Answer: Approximately 37.5%
Show your work in this space: P = 11 + 44ln 4
P = 11 + 44 × 0.602059
P = 11 + 26.490596
P = 37.490596
b) Use this model to determine in how many years it will take for the percent of spam to reach 85% provided that law enforcement regarding spammers does not change.
Answer:
Show your work in this space:
85=11+44 ln x
74=44 ln x (now am stuck)

Answer by scott8148(6628)

check your work for part a) __ it looks like you used log instead of ln


dividing by 44 __ 74/44=ln(x) __ taking antilog __ e^(74/44)=x